Trucks A (1000 kg) and B (2000 kg) are parked on a 10.0-m bridge. If A is

situated 5 m form the center of the bridge and B is parked 6.0 m from the other

end, how much force shall each bridge support hold? Use F1 for the support

near A and F2 for the support near B.


1
Expert's answer
2021-03-17T16:54:30-0400


Equilibrium of forces:


"F_1-mg-Mg+F_2=0"

Equilibrium of moments with left end as a pivot point:

"-Mg·4+F_2·6=0,\\\space\\
F_2=\frac{4Mg}{6}=13.1\text{ kN}."

From equilibrium of forces:


"F_1=g(m+M)-F_2=16.3\text{ kN}."